Find the length of each hypotenuse.
(Legs)^2+(legs)^2=hypotenuse^2
(39)^2+(24)^2=x^2

Answer by stanbon(75887)

(39)^2+(24)^2=x^2
---
x^2 = 2097
---
x = 45.793...
Cheers,
Stan H.